### v5

# Finding the more fundamental law of classical mechanics

Abstract
We have studied the relationship between the conservation of energy and Newton’s second law or equations of motion. Differently from recent works, we show that equations of motion can be obtained only in some special cases and their results are distinct in general. This distinctiveness makes it possible to find the more fundamental law experimentally.
1 Introduction
The fundamental laws constitute the bases of scientific understanding. The conservation of energy and Newton’s second law or equations of motion, which can be obtained from Lagrange equations, are two of the fundamental laws of classical mechanics. In general, these two laws are considered separate fundamental laws. This can be seen as natural, but there are some recent works considering relations between these, mostly trying to obtain equations of motion from the conservation of energy.
The history related to these laws is a long one and some parts are related to the ”vis viva controversy” including Newton’s second law and Leibniz’s usage of mass times velocity square [1, 2, 3]. Most of us had started to use Newton’s approach since Newton’s method was successful and Leibniz’s formulation had some problems. Later, the conservation of energy for mechanics is shown to be true by the work of scientists like Young, Coriolis, Helmholtz and many others [4, 5, 6]. There were also scientists trying to understand the relationship between Newton’s second law and the conservation of energy. D’Alembert, Lagrange and Carnot were aware that the conservation of energy can be seen as a consequence of Newton’s second law in some special cases [7]. These scientists took Newton’s second law as the basis and did not consider the other way around so do most of us. This is, probably, because of learning Newton’s second law prior to the conservation of energy and trying to understand it in terms of Newton’s second law.
Within centuries, we have learned lots of things, and Lagrangian formalism is used successfully nearly all over physics. If we consider this and see that it is possible to obtain the conservation of energy by using Newton’s second law, we may regard it as more fundamental than the conservation of energy. But, there are other works related to the topic.
Later, Hamilton developed another formalism to obtain equations describing the system, Hamilton equations, from Hamiltonian which is equal to energy ”if the equations defining the generalized coordinates don’t depend on time explicitly”, ”and if the forces are derivable from a conservative potential” [8]. Hamilton equations will provide equation sets in first-order differential equations for each momentum. If we consider the cases where Hamiltonian is equal to energy, we have a formalism giving equations describing the system from energy.
Recently, there are other approaches to this topic. Vinokurov gives a formalism, the principle of detailed energy conservation, to obtain equations of motion from a Lagrangian which is obtained in terms of energy by using the conservation of energy in his work [9]. Carlson writes energy in terms of generalized coordinates and momenta and then gives a formalism to get equations of motion by using the conservation of energy, and states that his formulation does not describe properly electromagnetic field and relativistic cases [10]. Lindgren presents a prescription, generalized energy method, to obtain equations of motion which is formed by using the relation that the work done by a force is equal to change in kinetic energy, which can be considered as the basis of the conservation of energy, and claims that equations of motion can be obtained except the cases including Coriolis force and gyro moments [11]. Zhou and Wang claim that they obtain equations of motion in different cases by using the conservation of energy [12]. Hanc and Taylor give a summary of Newtonian physics including Lagrange equations and the conservation of energy, and obtain the equation of motion for a one-dimensional motion from the conservation of energy while giving this summary [13]. On the other hand, Neuenschwander, Taylor and Tujela state that variation of energy is not a suitable quantity to obtain equations of motion [14].
As it can be seen there are various approaches to the problem and some scientists are trying to obtain equations of motion from the conservation of energy and some of them argue against it. We have mentioned above that one may regard Newton’s second law or equations of motion as more fundamental than the conservation of energy since one can show the conservation of energy by using it. On the other hand, if we approach the problem from the other side, we may reach a totally different conclusion. If we accept the results of obtaining equations of motion from the conservation of energy as correct and consider how wide the principle of the conservation of energy is, then we may regard it as a more fundamental law than Newton’s second law. This is a contradictory situation.
In this work, we will study the relationship between the conservation of energy and equations of motion without trying to obtain the same or different results. Though some authors claim that equations of motion can be obtained from the conservation of energy, we show that it is possible only in some cases. More importantly, the results show that there is a subtle difference between the conservation of energy and Newton’s second law or equations of motion. And, this difference is worth to further study. It can be used to determine experimentally the more fundamental law.
2 Theory
One can get equations of motion from Lagrange’s equations which are given by [8]
$\frac{d}{dt}\left(\frac{\partial\mathscr{L}}{\partial\dot{q}_{i}}\right)-\frac{\partial\mathscr{L}}{\partial q_{i}}=0,$ (1)
where $\mathscr{L}=T-U$ is Lagrangian, $T$ is the kinetic energy, $U$ is the potential energy, $t$ is the time, $q_{i}$ and $\dot{q}_{i}$ are generalized coordinates and velocities, respectively.
Now let us consider energy, $E(q_{i},\dot{q}_{i})=T+U$. For energy-conserved systems, its total time derivative is equal to zero. In this case, one can write
$\sum_{i}\left[\frac{\partial E}{\partial\dot{q}_{i}}\frac{d\dot{q}_{i}}{dt}+\frac{\partial E}{\partial q_{i}}\frac{dq_{i}}{dt}\right]=0.$ (2)
This equation is one of the starting points of previous works that are trying to relate the conservation of energy with equations of motion [9, 10]. It is hard to compare this equation with Lagrange equations in these forms. Forasmuch as we will consider scleronomic systems having velocity-independent potential energy, $U=U(q_{i})$. For scleronomic systems with many particles, kinetic energy can be written as $T=\frac{1}{2}\sum_{jk}M_{jk}\dot{q}_{j}\dot{q}_{k}$ where $M_{jk}=\sum_{i}m_{i}\frac{\partial\vec{r}_{i}}{\partial q_{j}}\cdot\frac{\partial\vec{r}_{i}}{\partial q_{k}}$, $m_{i}$ and $\vec{r}_{i}$ are the mass and position vector of $i$th particle, respectively [8]. Then, Lagrange equations for $i$th particle can be found as
$\sum_{k}M_{ik}\ddot{q}_{k}+\frac{1}{2}\sum_{jk}\frac{\partial M_{jk}}{\partial q_{i}}\dot{q}_{j}\dot{q}_{k}+\frac{\partial U}{\partial q_{i}}=0,$ (3)
and from Eq. (2), one can obtain
$\sum_{i}\dot{q}_{i}\left[\sum_{k}M_{ik}\ddot{q}_{k}+\frac{1}{2}\sum_{jk}\frac{\partial M_{jk}}{\partial q_{i}}\dot{q}_{j}\dot{q}_{k}+\frac{\partial U}{\partial q_{i}}\right]=0.$ (4)
This equation, obtained from the total time derivative of energy, has a summation over index $i$ and is distinct from equations of motion. Differently from equations of motion, the summation of all terms can be equal to zero.
From these two equations, similar to D’Alembert, Lagrange and Carnot, one may conclude that Lagrange equations are enough to show the conservation of energy for considered systems since the generalized velocity $\dot{q}_{i}$ is different from zero in general. This is also shown previously for some cases without considering scleronomic systems [15, 16]. But, the mentioned conclusion is the result of approaching the problem from only one perspective. And, this topic requires further consideration. We can widen our perspectives and ask a question to study further: Is the conservation of energy more fundamental than equations of motion or vice versa?
Now, we will try to find an answer to this question. If the motion is one-dimensional, then the summation over index $i$ in Eq. (4) drops. And, since $\dot{q}_{i}\neq 0$ in general, one obtains the same equation as Lagrange equations. Similar considerations are also valid for two or three-dimensional systems where by using conserved momenta, one can write Lagrangian and energy in a one-dimensional form. One can easily see these by studying central force motion and the heavy symmetric top.
Multi-dimensional cases, where the reduction is impossible, give more interesting results. As an example, one can consider the double pendulum whose kinetic and potential energies can be written as $T=\frac{1}{2}(m_{1}+m_{2})l_{1}^{2}\dot{\phi}_{1}^{2}+\frac{1}{2}m_{2}l_{2}^{2}\dot{\phi}_{2}^{2}+m_{2}l_{1}l_{2}\dot{\phi}_{1}\dot{\phi}_{2}\cos(\Delta\phi)$ and $U=(m_{1}+m_{2})gl_{1}(1-\cos\phi_{1})+m_{2}gl_{2}(1-\cos\phi_{2})$, respectively, where $m_{i}$ are masses of bobs, $l_{i}$ are lengths of pendula, $\phi_{i}$ are deviations from the equilibrium position and $\Delta\phi=\phi_{1}-\phi_{2}$ [17]. By using Eq. (2), one can obtain
$\dot{\phi}_{1}$ $\left[\frac{m_{1}+m_{2}}{m_{2}}\frac{l_{1}}{l_{2}}\ddot{\phi}_{1}+\ddot{\phi}_{2}\cos(\Delta\phi)+\frac{m_{1}+m_{2}}{m_{2}}\frac{g}{l_{2}}\sin\phi_{1}+\dot{\phi}_{2}^{2}\sin(\Delta\phi)\right]$ (5)
$+$ $\dot{\phi}_{2}\left[\frac{l_{2}}{l_{1}}\ddot{\phi}_{2}+\ddot{\phi}_{1}\cos(\Delta\phi)+\frac{g}{l_{1}}\sin\phi_{2}-\dot{\phi}_{1}^{2}\sin(\Delta\phi)\right]=0$
From Lagrange equations, equations of motion can be obtained as [18]
$\frac{m_{1}+m_{2}}{m_{2}}\frac{l_{1}}{l_{2}}\ddot{\phi}_{1}+\ddot{\phi}_{2}\cos(\Delta\phi)+\frac{m_{1}+m_{2}}{m_{2}}\frac{g}{l_{2}}\sin\phi_{1}+\dot{\phi}_{2}^{2}\sin(\Delta\phi)=0$ (6)
$\frac{l_{2}}{l_{1}}\ddot{\phi}_{2}+\ddot{\phi}_{1}\cos(\Delta\phi)+\frac{g}{l_{1}}\sin\phi_{2}-\dot{\phi}_{1}^{2}\sin(\Delta\phi)=0.$ (7)
The result of Eq. (5) and Lagrange equations are not equal to each other. However, it can be seen that the conservation of energy, i.e. $\frac{dE}{dt}=0$, can be obtained from equations of motion by using the fact that it is equal to $\dot{\phi}_{1}$ times Eq. (6) plus $\dot{\phi}_{2}$ times Eq. (7) and Eq.s (6) and Eq. (7) are equal to zero. On the other hand, though the conservation of energy can be obtained from equations of motion, the equation that is obtained from the conservation of energy, i.e. Eq. (5), is different and does not give equations of motion. There is a subtle point here which will be considered at the conclusion.
3 Conclusion
We have studied the relation between the conservation of energy and equations of motion for scleronomic cases with velocity-independent potentials. We have seen that one can get the same equations from the conservation of energy and equations of motion for one-dimensional cases and some multi-dimensional cases where energy can be written in one-dimensional form by using conserved momenta.
For multi-dimensional cases, whose energy cannot be written in one-dimensional form, the resultant equations are not the same. This difference shows that Lagrange equations and the conservation of energy do not designate the same results, in general. To see the difference, we have considered the double pendula as an example.
We have mentioned in the introduction that Hamiltonian is equal to the energy in some cases and Hamilton equations, a set of equations for each momenta, can be obtained from it. If we consider the double pendula example, there will be two sets of equations for each momenta. On the other hand, from the conservation of energy we obtain only a single equation. Then we can easily say that Hamilton equations and the equation obtained from the conservation of energy are different things.
There are some recent studies concerning the double pendula, and only some of these consider the conservation of energy [19, 20, 21]. However, none of these studies include any results considering any possible difference between Eq. (5) and Eq.s (6)& (7). Nevertheless, by studying experimentally the double pendula or a similar case, one can find an answer to the previously asked question: Is the conservation of energy more fundamental than equations of motion or vice versa?
An experiment can obey the equation obtained from the conservation of energy, i.e. Eq. (5), while not obeying equations of motion, i.e. Eq.s (6) and (7). If this case holds, then the conservation of energy is more fundamental, and equations of motion or Newton’s second law is a mere consequence of the conservation of energy for cases where either energy is one-dimensional or energy can be written in the one-dimensional form. A single experimental result is enough to say this.
On the other hand, an experiment can obey equations of motion, i.e. Eq.s (6) and (7), and the conservation of energy is satisfied as a natural result of this. In this case, a single result is not enough to say that the conservation of energy is the result of equations of motion, since it does not cover all possibilities.